OpenStudy (anonymous):

Someone please help its my last question. A car is traveling at a speed of miles per hour. What is the car's speed in miles per minute? How many miles will the car travel in minutes? Do not round your answers.

OpenStudy (anonymous):

How many minutes in an hour?

OpenStudy (anonymous):

60

OpenStudy (anonymous):

If I drive 1 mile per minute, how many miles per hour did I drive?

OpenStudy (anonymous):

60?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

So if you drive 60 miles per hour, it is only 1 mile per minute. So to convert, it seems you just divide by 60.

OpenStudy (anonymous):

\[ \frac{miles}{\cancel {hour}}\times \frac{\cancel {hours}}{60\ minute} = \frac{miles }{60\ minute} \]

OpenStudy (anonymous):

Oh wait im so sorry its supposed to say 93 miles per hour and 10 minutes

OpenStudy (anonymous):

Okay so first convert to miles / minute like it asks.

OpenStudy (anonymous):

\[ \frac{93\ miles}{hour}=\frac{93\ miles}{60\ minute} \]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

@wio ?

OpenStudy (anonymous):

Sorry I'm a bit tired. I hope someone else can help.

OpenStudy (anonymous):

Oh its okay :/ thanks thought